One urn contains 3 read and 7 white marbles. A second urn contains 2 red and 8 white marbles. One urn is selected and a ball drawn. The probability of selecting the first urn is 0.4 and the probability of selecting the second is 0.6. If a white ball is drawn, find the probability it came from the second urn.
Answer by stanbon(75887)

One urn contains 3 read and 7 white marbles. A second urn contains 2 red and 8 white marbles. One urn is selected and a ball drawn. The probability of selecting the first urn is 0.4 and the probability of selecting the second is 0.6. If a white ball is drawn, find the probability it came from the second urn.
---------
P(2nd urn|white) = P(2nd and white)/P(w)
-------------
But P(2nd and white) = 8/20 = 2/5 = 0.4
---------
And P(w) = P(w |1st urn)+P(w |2nd urn)
=P(w AND 1st)/P(1st)+P(w AND 2nd)/P(2nd)
= (7/20)/(8/20 + (8/20)/(12/20)
= (7/8) + (2/3)
= 37/24
---------------
Final: P(2nd urn |white) = 0.4/(37/24) = 0.259...
===========
If you have another answer I would like to see it.
Why? Probability problems are always tricky.
=============
Cheers,
Stan H